Water flows at 3m/s through a pipe of diameter0.1m and strikes a vertical plate near the outlet of the pipe. If the stream of water strikes the plate normally,calculate the force exerted in the plate( density of water is 1000kg/m
1
Expert's answer
2020-06-15T10:34:38-0400
"F=pA=(0.5\rho v^2)(0.25\pi d^2)"

"F=(0.5(1000)(3)^2)(0.25\pi (0.1)^2)=35\ N"